OpenStudy (anonymous):

how would it work for solving |7x|−12≤9

OpenStudy (anonymous):

start by adding 12 to get \(|7x|\leq 21\)

OpenStudy (anonymous):

then solve one compound inequality \[-21\leq 7x\leq 21\] in one step by dividing all by \(7\)

OpenStudy (anonymous):

so i end up with x=3 and x=-3, right?